OpenStudy (anonymous):

A 75-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. Find the horizontal and vertical distances spanned by this brace. So far I have: Vertical distance=x horizontal distance=x+15 Pythagorean Theorem a^2+b^2=c^2 x^2+(x+15)^2=75^2 x^2+x^2+30x+225=5625 0=2x^2+30x-5400 a=2 b=30 c=-5400

OpenStudy (anonymous):

were did you get x+15?

OpenStudy (anonymous):

cant solve the problem coz its lack of given variable

OpenStudy (anonymous):

Some one else on here was helping me with the problem. I do not understand their way of doing it. This week it was factoring and the Pythagorean Theorem we did that too.

OpenStudy (anonymous):

Could we just solve it using the Pythagorean Theorem? I can understand to use that, otherwise it makes no sense to use something else.

OpenStudy (anonymous):

i guess you can solve also using Pythagorean Theorem..but i was thinking of x+15 coz its not specified in the problem

OpenStudy (anonymous):

Well, lets just do the problem. without any of that other stuff I have so far! It wasn't making sense to me, lets start over.

OpenStudy (anonymous):

but for me theres a lack of given variable:-)

OpenStudy (anonymous):

I typed it just like it was. Maybe I could retype it. We have to find the variable?

OpenStudy (anonymous):

What chris is implying is you either need a side measurement or an angle.

OpenStudy (anonymous):

If it's a 45 degree angle then the length and height would equal for example.

OpenStudy (anonymous):

okay..kindly review the problem.maybe there is lacking.thanks

OpenStudy (anonymous):

oh, but then at the end it just says find the horizontal and vertical distances spanned by this brace. So thats what we have to figure out

OpenStudy (anonymous):

so if this the case that you gave us this given Vertical distance=x horizontal distance=x+15.i`ll try to solve if this is a correct answer. Use Pythgoras to find the missing sides of your triangle: 75^2 = x^2 + (x+15)^2 x^2 + x^2+30x+225 = 75^2 2x^2 + 30x - 5400 = 0 x^2 + 15x -2700 = 0 (x-45)(x+60) = 0 Positive solution: x = 45 ft (vertical distance of the brace) x+15 = 60 ft (horizontal distance of the brace)

OpenStudy (anonymous):

A 75-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 15 ft longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace. I was just thinking if a bridge connects a support of the center of the bridge to a side support on the bridge. Would this be a 90 degree angle?

OpenStudy (anonymous):

Yes it would. See nicks solution. The part that was missing was the 15ft longer part.

OpenStudy (anonymous):

Err chriss solution.

OpenStudy (anonymous):

ahhh this sentence was lacking "The horizontal distance that it spans is 15 ft longer than the height that it reaches on the side of the bridge"

OpenStudy (anonymous):

thats what im implying for coz i was wondering where did u get this x+15 lately..:-)

OpenStudy (anonymous):

whats err by the way? :-)

OpenStudy (anonymous):

Well when you look at it, the beginning of the last two sentences you can see in the paragraph is the horizontal. So sorry! I had I feeling I was talking to my mom and not making sure it was all there. Good thing you have good eyes to notice that.. Thank you for your help guys, heres a medal!

OpenStudy (anonymous):

glad to help you..your welcome and thanks